OpenStudy (anonymous):

Ok i started my job yesterday and i get paid 9.14 an hour every two weeks i work 5 hours a day for 4 days how much would my paycheck be 2 weeks from now

OpenStudy (anonymous):

no. of days=2*4=8 no. of hours==8*5=40 amount of check=40*9.14=?

OpenStudy (anonymous):

365.00 not considering Taxes and etc to be deducted